Register Login Password forgotten?

Station History Log - Graz, Austria (7839)

SOD Year Doy Toy System Description
78393402 2019 070 15:00 99 Pulse Distribution Box / Power Supply repaired, cables changed, affects cal constant
78393402 2015 124 08:00 5.02 Laser Beam path on laser table changed; causes CAL constant change only
78393402 2015 023 00:00 99 Mount Main dichroic changed to IQOQI 2 (reflects 800 nm, transmits everything else)
78393402 2015 023 00:00 99 Mount DetPack New M1/Dichroic (EO86-400): Rmax 532 / Tmax 650-1200 nm
78393402 2015 023 00:00 99 Mount Temp stab for C-SPAD housing directly coupled to temp controlled air flow
78393402 2014 289 00:00 99 Mount Complete mechanical rebuild of detetction package; but same optical / electronical units
78393402 2014 181 05:00 99 RANGE Program updated V7.31: Format#5 for returns; increased for >99999 SoD
78393402 2014 016 11:00 99 RANGE Program updated V7.30: FPGA RG FIFO reduced, part of future RGs stored now in PC
78393402 2013 181 00:00 99 CPF_WG 2x Iteration changed to 4x Iteration, for horizontal HEO residuals
78393402 2013 171 15:00 99 Laser Room at 15:00 UTC; Temperature reduced by 2C to 20C (mount Electronics)
78393402 2013 163 00:00 99 Laser Polarization fixed again; Unit disconnected, software still sends commands
78393402 2013 074 00:00 8.01 CAL CAL Program extended for different lasers; new setup files; calfile format adapted
78393402 2012 324 00:00 5.02 Laser Lambda/4 removed; Lambda/2 RC unit re-installed; linear polarization; restart Glonass Retro Tests
78393402 2012 289 00:00 99 Coude Coude mirrors: Cleaned (with Polymer); new airtight path; aligned
78393402 2012 269 00:00 4.01 Mount Az Poti included
78393402 2012 254 00:00 99 Range NPts_AG: Valid points only if AG is on
78393402 2012 208 00:00 99 PDM PDM 2012 installed: Improved NIM; 10kHz ready
78393402 2012 083 00:00 5.02 Laser Lambda/4 plate installed; changes linear into circular poarization; starts with LAS08309
78393402 2011 286 00:00 99 ET 10 kHz System using Graz RIGA ET again
78393402 2011 262 00:00 99 PDM Old (Version red/blue/green front) Pulse Distribution Unit installed: better jitter behaviour then version above; new CAL value
78393402 2011 259 00:00 99 PDM Old (Version red/blue front) Pulse Distribution Unit reinstalled: 10 kHz Version - problems with NIM pulse; new CAL value
78393402 2011 257 00:00 99 ET 10 kHz System using the San Fernando RIGA ET
78393402 2011 255 00:00 99 Start Start Pulse Detector Jan1 installed
78393402 2011 250 00:00 5.02 Laser Laser + Start pulse detector new adjustment
78393402 2011 243 00:00 99 PDM Pulse Distribution Module: Trigger Level increased from 40 mV to 80 mV
78393402 2011 203 00:00 99 TIME HP GPS: Antenna delay was ZERO; new Antenna Delay 153 ns programmed
78393402 2011 180 00:00 99 RANGE7.21 Histogram plots: Includes nonlinear noise, higher resolution etc.
78393402 2011 157 00:00 99 Filter Blocking filter replaced: OD6 / 93% Transmission / 535 nm CWL, 43 nm BW
78393402 2011 105 00:00 99 MCU Error in DE_NANO FPGA removed: Daylight filter now removed again during night
78393402 2011 105 00:00 99 Filter Blocked 0.3 nm (35%) filter replaced with unblocked 0.3 nm (48%); CAL quote from 21% to 38%;TFilter: 22 ps
78393402 2011 100 00:00 99 Filter Blocking filter (90%, 10 nm) on top of wavelength filter installed
78393402 2011 096 00:00 99 MCU DE_NANO FPGA on mount installed; daylight Filter unintentionally not removed during night (but TFilter applied)
78393402 2011 063 00:00 4.01 Mount New Dichroic / KV Optical installed; Mount Subroutine: Fraction from 0.5 to 0.2 changed; dramatic improvement
78393402 2011 027 00:00 5.02 Laser New Startpulse Detector Version 08-2010 (PDM Version)
78393402 2011 026 00:00 99 RANGE 4-Flag-Format
78393402 2011 019 00:00 99 Laser Laser box placed on new position; 2 telescopes and old stuff removed; new CAL value
78393402 2011 018 00:00 99 RANGE Write to Win_PC: PC104/COM1; Sun Angle; Send_Transp;
78393402 2011 012 00:00 99 PDM New Pulse Distribution Unit: 10 kHz Version, with disable / enable bits etc.; new CAL value
78393402 2011 011 00:00 5.02 Laser New crystal in Regen Amp; new adjustment;less current in Regen URDM; 5 kV Pockels
78393402 2010 280 00:00 5.02 Laser New adjustment, new Pockels Cell, 5 kV, 400 uJ
78393402 2010 264 06:00 4.01 FILTER Tubus Temperature Control failure detected; fuse resetted
78393402 2010 253 00:00 5.02 Laser Laser Adjustment Problems with new Pockels Cell started;
78393402 2010 245 00:00 5.02 Laser New Pockels Cell installed
78393402 2010 200 00:00 5.02 Laser Pockels Cell complete readjustment
78393402 2010 181 00:00 5.02 Laser Pockels Cell: Delay + Pulse Width adjusted; less trailing pulses
78393402 2010 181 00:00 4.01 Mount Elevation: Modules Encoder; Course; Course/Fine replaced with RGO modules
78393402 2010 179 00:00 99 ISA FPGA Overlap Check Time: From 60 us increased to 80 us
78393402 2010 159 00:00 99 WinClean Version 1.3: NP CRD files routinely stored on DGFI ftp side, parallel to old NPs
78393402 2010 126 00:00 99 WinClean Version 1.3: C:Laset2010NPNP126 ...; first CRD NPs (tests only) started
78393402 2010 117 00:00 99 CAL CAL opening reduction removed; standard again
78393402 2010 110 00:00 6.01.01 SPAD SPAD aligned, using CAL target; CAL opening reduced to 1/3
78393402 2010 105 00:00 99 FOV FOV Servo with toothed wheel,for ZERO slackness, no backlash etc.
78393402 2010 036 00:00 99 RANGE ET1_Ep + TFilter => ET_TI + TFilter
78393402 2010 036 00:00 99 BLITS Filter OUT for BLITS Night Passes
78393402 2010 035 00:00 99 RANGE Day/Night Switch (RED LED) instead of LCD; Switch 1/2/4/Speed obsolete (or done via program);
78393402 2009 320 00:00 5.02 LASER Adjusted:No Pre-train, a bit more post-tracks
78393402 2009 300 00:00 5.02 LASER Oscillator realigned; crazy tracks before (weekend + day 300)
78393402 2009 295 00:00 5.02 LASER REGEN + Post readjusted; Indium Seal blocked part of beam; old Regen Pump Diodes reinstalled
78393402 2009 281 00:00 99 FILTER OUT for BLS / Night Passes; did not work !!!
78393402 2009 274 00:00 99 RG Kabel fuer RG geaendert; RANGE: Del0 von 135 ns auf 160 ns
78393402 2009 273 00:00 99 RANGE CO-Filter: IDet_Limit fuer HEOs von 100ps auf 300 ps geaendert
78393402 2009 264 00:00 4.01 FILTER OUT for Night/HEO/AUTO: 15 ps added for missing filter glas
78393402 2009 246 00:00 99 RANGE Mount_Loop: 30 ms/100 ms Update Rates for HEO/LEO Satellites
78393402 2009 223 00:00 4.01 ISIT RCA ISIT (US system) defect, JAI ISIT installed (European System); no Seeing values derived
78393402 2009 195 00:00 5.02 LASER Pockels Cell readjusted; Leading edge of rets back to normal; energy not stable; 1-s CHA intervals ?
78393402 2009 194 00:00 5.02 LASER REGEN: Pump Diodes replaced; Pockels Cell adjusted; increased jitter @ strong returns
78393402 2009 188 06:00 99 FPGA Times Before: 165.66-29.95=135.71 ns; After: 145.23-25.33=119.90 ns
78393402 2009 188 00:00 99 ISA FPGA RGG: Min/Max Delay corrected; before: up to 20 ns drift / 100 ns interpolation (analog chip)
78393402 2009 183 00:00 99 RANGE Mount_Loop: From 30 ms to 100 ms update intervals changed
78393402 2009 182 00:00 99 Station Thunderstorm / Flash: Serial Interface to Dome / Dome Frequency Converter damaged
78393402 2009 105 00:00 99 Station Power failure for some hours
78393402 2009 061 00:00 99 PC CPU Card defect; exchanged ....
78393402 2009 043 00:00 99 OCU OCU and MCU operational
78393402 2009 015 00:00 99 Format Epochs: From F14.8 to F18.12 (10 ns to 1 ps; only 10-100 ps acc); WinClean, Range changed
78393402 2009 014 00:00 99 SatDef C:LasetDefSatDef replaces all other sats/sat.cpf etc. files; all progs changed
78393402 2009 014 00:00 99 Dome Observer Cabin: New Table / PCs disconnected / new connections etc.
78393402 2008 342 00:00 99 Dome Station up; new dome / RS232 control implemented in Range
78393402 2008 311 00:00 99 Dome Station down; Old Dome Removed; all cables OFF etc
78393402 2008 198 00:00 99 Transmit Testing New Transmit Telescope (20 cm lens); no data sent
78393402 2008 197 00:00 99 CPF Predicts now with New PC/XP, also the ftp transfer / Zipped
78393402 2008 178 00:00 99 KG PC New PC/XP; Softcal etc. LF9045
78393402 2008 161 00:00 99 MET PC New PC / XP
78393402 2008 065 00:00 99 WinClean AJI: LE /20 mm/ 5 mm RMS started
78393402 2008 063 00:00 5.02 Start Start Pulse Detector adjusted: From 2.5 V to 3.5 V
78393402 2008 059 00:00 5.02 CAL Start Pulse Detector: BNC/SM/BNC changed to SMA only; 200 ps jump in CAL value
78393402 2008 047 00:00 99 CLEAN Error Writes to Statistics corrected: Pts/NP; Quote were wrong ...
78393402 2008 043 00:00 6.01.01 SPAD Realigned; Laser realigned; now much better results
78393402 2008 036 00:00 99 WinClean LA1/LA2: Leading Edge /20 mm/ started (5mm RMS); AJI: LE /7 mm RMS / started
78393402 2008 023 00:00 99 CLEAN New WinClean Program operational; mainly doing manual, but also few AutoCleans
78393402 2007 263 00:00 12.03.01 MET External analog MET sensors transferred into Laser Room
78393402 2007 262 00:00 6.01.01 ET ET_Sync; maybe power failure; thunderstorm
78393402 2007 261 00:00 99 GPS HP GPS Antenna: Back to (roughly) previous location; restart
78393402 2007 235 06:00 99 Mount Power Supply / Sun Shutter etc. failure / worked again ...
78393402 2007 234 00:00 99 Mount Power Supply Failure; repaired; fuse
78393402 2007 234 00:00 99 Mount Several Failures due to frequent OFF/ON (observatory)
78393402 2007 211 00:00 10.01 CLEAN Results: Noise points showed up to now an offset of +1.3 m (SemiTrain Distance ...); corrected
78393402 2007 205 00:00 99 GPS HP GPS Antenna: New Location: Roof Middle; + 403 cm higher, same cable length, switched OFF/ON
78393402 2007 177 00:00 99 LAS PC All relevant Programs with COM_Select (NN): TWO Writes to CAS-81 instead of ONE; works again
78393402 2007 170 00:00 99 LAS PC CF Card replaced by standard disk again
78393402 2007 169 00:00 99 LAS PC CF Card 2 GB used instead of Hard Disk; all systems on CF card
78393402 2006 258 00:00 99 Mount Cabl RG58/SMA Stop1 from SPAD to Base Plate installed
78393402 2006 250 00:00 99 Mount Cabl New RG58/BNC cable from SPAD to Contraves base plate (flexible core) instead of RG223 (rigid core)
78393402 2006 249 00:00 99 Mount Cabl jumps; tests with fiber; due to cable/AZ/crosstalk? cable ties changed; better, but not solved
78393402 2006 241 09:00 99 Mount Cabl 9:00 Same jumps again(60 ps): Contact spray on Az Base SMA Connectors, Stop4;
78393402 2006 241 00:00 99 Int PCB Temp Control reconnected; +- few ps again
78393402 2006 240 12:10 99 Mount Cabl 12:10 SMA Cable for Stop Pulse: Changed from Stop1 to Stop3 SMA Cable
78393402 2006 237 16:00 99 Mount Cabl 16:00 4 SMA cables RG223 at yoke from outside touched, big cal variations; after that okay ...
78393402 2006 236 14:30 6.01.01 C-SPAD 14:30 => Internal PCB Temp Control (Heating) disconnected. T decreases => Cal decreases
78393402 2006 233 13:00 6.01.01 C-SPAD From 233 - 234 / 13:00 UTC: CAL var +-40 ps, Temp Sensor SPAD Housing "obigfoin"
78393402 2006 212 00:00 5.02 kHz Laser Start Pulse aligned; 3 days before bigger variations; temp sensor was not on C-SPAD anymore
78393402 2006 191 00:00 99 CPF Pred CPF predictions; fully automatic preds since day 198
78393402 2006 053 00:00 5.02 kHz Laser Laser readjusted, REGEN Pump increased by 0.2 A; Osc. decreased by 100 mA
78393402 2006 045 00:00 99 kHz Laser Start Puls Detector Position readjusted; from +2V to +3V pulses
78393402 2006 016 00:00 99 LAS PC Switch to Industry PC/3.2 GHz, with IEC-Bus/LPM16/DIO96 Cards; No ISA/Altera
78393402 2005 314 00:00 99 MET PC HD Failure during rearrangements; on evening everything restarted again
78393402 2005 312 00:00 5.02 kHz Laser Oscillator Adjust.; complete adjust; 1050 to 900 mA Osc; 375 Units
78393402 2005 311 00:00 99 Laser Room 2 Racks removed; all devices in Rack#3;(Rack4: Mount); IEC-Bus: Needs 5328
78393402 2005 297 00:00 5.02 kHz Laser Osci Dioden same; nachjust no change; Start Pulse: from 2.7 => 3.0 V
78393402 2005 290 06:00 5.02 kHz Laser Oscillator Pump Diodes: +50 mA; RMS slightly degraded then
78393402 2004 212 00:00 99 CALK After Laser Readj.(208): 5 Round trips instead of 6-7, RG was slightly too late; corrected now
78393402 2004 210 00:00 12.01.01 MET3A Recalibrated; installed again; Before: 0.290 hPa diff; now: 0.188 hPa diff
78393402 2004 208 00:00 5.02 kHz Laser Schmidt: Complete new alignment; new crystal postamp; C-SPAD aligned; RMS CAL from 14 now 20 ps
78393402 2004 189 00:00 12.01.02 MET3A Removed for Recalibration; No MET-PC Records; Using Paroscientific + external sensors
78393402 2004 093 00:00 6.01.01 SPAD PS New Spad Power Supply installed; PS3-10; uncompensated Out now via RG224 Stop #2
78393402 2004 058 00:00 6.01.01 TWC TWC new adjustment; from 110 ps to <10 ps; Power Supply Poti Problem
78393402 2004 028 00:00 5.02 kHz Laser Mr. Frei (HQ) fixed the crystal of the RegenAMP and the 2.AMP;
78393402 2004 028 00:00 5.02 New Alignment (RM1, RM2 etc.; Start Pulse Sensor); from 840 AU up to 940 AU;
78393402 2004 021 00:00 5.02 kHz Laser Spad Unit 0406 / Test Ch. Timewalk new adjusted PS: -6V to -6,450V
78393402 2004 021 00:00 5.02 kHz Laser SESAM: from 1000mA to 700mA no change in energy, timewalk and cal-time
78393402 2004 021 00:00 5.02 kHz Laser New Alignment (RM1, RM2 etc.); from 810 AU up to 910 AU;
78393402 2004 016 00:00 5.02 kHz Laser SESAM: from 700mA to back 1000mA ?SESAM, pulse length / time walk??
78393402 2004 015 00:00 5.02 kHz Laser pulse distribution box reinstalled; Triggervalues: Start=400mV, Stop=150mV
78393402 2004 013 00:00 5.02 kHz Laser SESAM: from 1000mA to 700mA no change in energy and cal-time
78393402 2004 013 00:00 5.02 kHz Laser max energy on start pulse detector (lower timewalk for energy changes, warmup)
78393402 2004 013 00:00 5.02 kHz Laser pulse distribution box removed (defect)
78393402 2004 007 00:00 12.03.02 kHz Laser New Alignment (RM1, RM2 etc.); from 760 AU up to 890 AU;
78393402 2003 345 00:00 9.02.01 GPS HP GPS PS fan 5V Supply only (Oven could not stabilize)
78393402 2003 331 00:00 6.01.01 TWC SPAD Time Walk Compensation: Adjusted for 10 ps kHz Laser
78393402 2003 331 00:00 9.02.01 GPS HP GPS new power supply installed (incl. 12V fan)
78393402 2003 324 00:00 5.02 kHz Laser New Alignment (RM1, RM2 etc.); from 0.25 mJ up to 0.40 mJ again;
78393402 2003 195 00:00 99 RGG RGG2, with FIFOs (512); Range 5.40; runs up to 1 kHz (NO Plot_Resids, but WITH MOUNT etc.)
78393402 2003 170 00:00 99 Computer SPECTRA 2.4 GHz ISA Pentium-4 Board; Running Range 5.31; Free running, MJS RGG
78393402 2003 015 06:00 12.01.01 PARO ParoScientific Digiquartz: New Calibration, running parallel, visual comparison
78393402 2002 344 00:00 12.01.01 MET3A RANGE/CAL/MET: Using MET3A ONLY + PC_Met; ParoScientific DigiQuartz => Calibration
78393402 2002 331 00:00 99 RANGE RANGE 5.31: Free running, MJS-RGG; 2xSR620 read/Store; CAL: 10-Hz-Version
78393402 2002 330 00:00 99 DOME New floor at Dome; all cables dis-/re-connected
78393402 2002 330 00:00 6.01.01 C-SPAD TWC new adjustment; -98 ps @ 1000 Photons before(!!!); now +3 ps; see SPAD Unit 406
78393402 2002 322 00:00 99 RANGX AutoSwitch for CAL (DG535/OLD); Laser Control Bit at F3 instead of Energy Control
78393402 2002 316 00:00 99 RANGX All Passes now with Free Running Range/RGG/10Hz ALL (5.30); CAL with DG535/OLD
78393402 2002 255 00:00 99 PENTIUM Pentium 200 MHz replaces PC486/66MHz
78393402 2002 183 00:00 99 CABLES 4 new cables RG223 (SMA1-SMA4, for STOP1-STOP4); new CAL Constant
78393402 2002 182 00:00 99 CABLES Test Channel / Stop 1: New Cable
78393402 2002 170 00:00 99 RANGE SPAD Gate Safety: From 120 nm to 51 ns (as previously)
78393402 2002 163 00:00 6.01.01 Filter New Filter (Andover 2002, 0.3 nm, 30%) installed
78393402 2002 162 00:00 5.01 Laser Nd:YAG, 532 nm, 35 ps, SemiTrain, 30 mJ
78393402 2002 162 00:00 6.01.01 Detector C-SPAD, 200 um, cooled to -60 degrees
78393402 2002 162 00:00 6.01.01 Counter ET; Clock Module, 1 Start, 1 Stop Module;
78393402 2002 162 00:00 99 Computer PC-486, 1xHPIB, Analog/Dig I/O;
78393402 2002 162 00:00 99 Range Gate Via HP5359 + DG535
78393402 2002 162 00:00 99 Software V 5.3, Tandems, OMO/GEO Mount system, Lagrange
78393402 2002 162 00:00 9.01.01 Timing 1Hz / 10 MHz via HP GPS
78393402 2002 162 00:00 12.02.01 MET ParoScientific #1; Home made temp/Hum sensors
78393402 2002 137 00:00 6.01.01 ET Module 3 disconnected
78393402 2001 094 00:00 5.01 LASER Oscillator + Amplifiers + Power GROUNDED
78393402 2001 031 00:00 6.01.01 ET From CH1 to Ch6
78393402 2001 030 00:00 9.01.01 HP5370A Fan broken; changed from 110 V AC to 220 V AC
78393402 2001 029 06:00 9.02.01 GPS New Power Supply installed for HP GPS
78393402 2001 025 00:00 6.01.01 ET Back to Ch1; new PCB ECL to TTL
78393402 2001 019 00:00 6.01.01 ET ET: From Ch1 to Ch3
78393402 2001 005 00:00 99 POWER Electrical Power failure; 3 h no power; ET synced
78393402 2001 005 00:00 9.02.01 GPS HP GPS Power Failure; external Supply
78393402 2001 001 00:00 6.01.01 ET New Synced
78393402 2000 214 00:00 6.01.01 ET ET ONLY; no counters; HP remains for Checks, sync;
78393402 2000 018 00:00 99 COUNTER Jitter in Counter Diffs; Change of 10 MHz Distr.; (KF SLR Timing)
78393402 2000 010 00:00 9.01.01 TIME UTC(TUG) disconnected by D. Kirchner; HP GPS now used
78393402 1999 328 00:00 5.01 START New PCB Start Pulse Detekt. + PrePulseDetekt.
78393402 1999 307 00:00 5.01 START New Position; new Cabeling of Start Pulse Detektor SFH229
78393402 1999 293 00:00 6.01.01 SPAD Heating Unit MOunt Detection Package
78393402 1999 291 00:00 99 CABLES New Cabeling f New Comparator Unit
78393402 1999 288 00:00 99 COMPAR New Comparator Unit; Delay DG535=100 ns;
78393402 1999 278 00:00 99 COUNTER COMP1/COMP2 disconnected from computer/OUT/IN/RECON.
78393402 1999 270 00:00 5.01 START New Start Pulse Detector SFH229
78393402 1999 253 00:00 99 Air Cond Sensor moved to Air Exit
78393402 1999 062 00:00 5.01 DYE Complete new Dye
78393402 1999 061 00:00 5.01 DYE CELL Flowing Dye Cell; Semitrain Distance 5752 ps
78393402 1999 054 00:00 5.01 SEMITRAIN Distance now: 8757 ps (before: 8759 ps)
78393402 1998 349 00:00 99 COUNTER CAL/RANGE: HP5370 as reference again
78393402 1997 364 00:00 99 PREPY RG: from 1 to 6 us suggested start value
78393402 1997 343 00:00 99 CABLES OUT1-Test: BNC4 ==> BNC8
78393402 1997 343 00:00 99 Air Cond RH_Int: From 2% to 22% (Poti ZERO)
78393402 1997 168 00:00 6.01.01 SPAD TWC new adjustment
78393402 1997 132 00:00 6.01.01 SPAD 29.5 ==> 29.8 V
78393402 1997 126 06:00 99 COUNTER SR620#3 added again
78393402 1997 114 00:00 99 COUNTER SR620#3 removed; repair 50 Ohm Stop
78393402 1997 112 00:00 99 FORMAT CLEAN: Format 1 implemented
78393402 1997 073 00:00 99 PC PentiumPro 200 MHz at Office, for Cleans etc.
78393402 1997 059 00:00 99 FORMAT Range/Clean: Format 1: Color/NTrack/Epoch/4xCounter
78393402 1997 057 00:00 9.01.01 TIME Sync Divider: Changes in 1 kHz Epoch Timing
78393402 1997 042 00:00 6.01.01 SPAD Daylight filter added
78393402 1997 041 00:00 6.01.01 SPAD dismounted, dried, mounted, adjusted; no daylight filter
78393402 1997 034 00:00 99 CAL/RANGE HP/SR620A results exchanged; SR620A now used as reference
78393402 1997 030 00:00 99 TIME Back to UTC(TUG) again ...
78393402 1997 023 00:00 99 TIME HP GPS DEMO used
78393402 1996 355 00:00 6.01.01 SPAD STOP SPAD STOP Somp Box Trigger: 250 ==> 100 mV (25 ==> 20 ps)
78393402 1996 334 00:00 6.01.01 Trigger Pre-Trigger-Unit adjusted, CAL now okay again
78393402 1996 296 00:00 99 Counters SR620C removed
78393402 1996 271 00:00 10.01 NP Min 5 Pts/NP minimum, otherwise NP not delivered
78393402 1996 271 00:00 6.01.01 TWC New, excellent adjustment
78393402 1996 269 00:00 99 Counters No longer f-daisy-chaining of counters; separate 10-MHz Supplied
78393402 1996 260 00:00 6.01.01 TWC Jitter after 6 days interruption
78393402 1996 254 00:00 6.01 SR620C Added; all 4 counters in use;
78393402 1996 253 00:00 99 PC IEC Bus cables changed to PC-Cables
78393402 1996 215 00:00 6.01 SPAD HP3631 Power Supply for SPAD; Remot Control via SPAD ON
78393402 1996 198 00:00 6.01.01 Trigger New Trigger Devices / MultiOutputs for Start/Stop
78393402 1996 170 00:00 99 Air Cond Air Temperature Control for hp only; SR620A/B: orig. fan setting
78393402 1996 144 00:00 12.01.02 ParoScientific Sensor: +21 cm higher; software corrected
78393402 1996 143 00:00 99 Temp Controlled Air for 3 Counters Cooling
78393402 1996 133 00:00 99 New Cabling of Mount/SPAD/Controls
78393402 1996 107 06:00 8.01 Target03 used, beginning 06:00 UTC
78393402 1996 060 00:00 4.01 Daylight filter installed again in Standard Channel
78393402 1996 059 00:00 6.01.01 SPAD adjusted again
78393402 1996 058 00:00 6.01.01 SPAD+TWC: From Test into Standard Channel
78393402 1996 054 00:00 8.01 Softcal, using Calfile
78393402 1996 046 00:00 8.01 CAL: Actual Met Vals used; CalFile Option available
78393402 1996 030 00:00 6.01.01 Clean routinely with 3 Counters
78393402 1996 025 00:00 6.01.01 CAL/RANGE: 3 Counters used, Results stored; not yet in CLEAN
78393402 1995 269 00:00 99 Sat_Char*3: New Pass Names
78393402 1995 265 00:00 12.01.01 ParoScientific Pressure Sensor installed
78393402 1995 261 00:00 6.01.01 SPAD VOltage set via PC Program
78393402 1995 238 00:00 99 USNO correction now also applied in PC
78393402 1995 234 00:00 99 First pass with PC
78393402 1995 016 00:00 6.01.01 HP5370A: 0.25/0.21 V Trigger Threshold (Before: 0.25/0.17V)